The original cost of the shirt was $30.00. Two weeks later the cost of the shirt was $24.00. What percent was it discounted?
Answer by ewatrrr(24785)


Hi
The original cost of the shirt was $30.00, discounted to $24
What percent was it discounted? (30-24)/30 = 6/30 = .20 or 20% discount
CHECKING our Answer***
$30 - .20*$30 = $30 - $6 = $24
OR 'paying 80%
.80*$30 = $24